OpenStudy (anonymous):

last month,a factory produced 800 television sets. This month, the same factory produced 1,064 t.v. sets the increase in production from last month to this month is?

OpenStudy (anonymous):

This is a simple subtraction problem. 1064 - 800 = 264

OpenStudy (anonymous):

If you're looking for a percent, then 264 / 800 = 33/100, so a 33% increase.